OpenStudy (laronjaxon):

Explain, in complete sentences, which method you would use to solve the following system of equations and why you chose that method. Show your work and provide the solution to the system. 4x + 3y = 2 3x + 2y = 1

OpenStudy (anonymous):

i would use multiplicativity,..

OpenStudy (shubhamsrg):

i'd use hit and trial cause thats' what makes me lazy :D

OpenStudy (anonymous):

substitution method.

OpenStudy (anonymous):

multiply the complete 1st equation by 2, so we get 8x+6y=4 multiply the complete 2nd equation by3, so we get 9x+6y=3 now its a cakewalk:) subtract both equations, u get (x,y)=(-1,12)

OpenStudy (anonymous):

its sleek for smaller equations, also it gave x with 1 as its coffecient, thts y i used it

OpenStudy (shubhamsrg):

you mean -1,2 right ? @dg123

OpenStudy (anonymous):

sorry sorry yes (-1,2) only

OpenStudy (shubhamsrg):

:D

OpenStudy (anonymous):

Substract (2) from (1) => x + y = 1 Multiply (2) by 2 => 2x + y = 0 Solve :)

OpenStudy (laronjaxon):

Thank you everyone

OpenStudy (laronjaxon):

@dj123 how come you chose that method?

OpenStudy (maheshmeghwal9):

@dg123

OpenStudy (laronjaxon):

thank u

OpenStudy (maheshmeghwal9):

welcome

OpenStudy (anonymous):

dis method was relatibely easy, nd the coffecicients strikes a bell, i got the feeling that dis would be the best way to do

OpenStudy (anonymous):

ts sleek for smaller equations, also it gave x with 1 as its coffecient(while subtracting), thts y i used it ,